OpenStudy (anonymous):

Solving Quadratic Functions 36x^2=100 36x-100=0 9x^2-25 root 9= 3 root 25 = 5 Factorized: (3x+5)(3x-5) My Ans: -5/3, 5/3 Is this correct?

OpenStudy (anonymous):

x^2 = 100/36 (divide by 36 on both sides) x^2 = 25/9 (reduce fraction) x = sqrt(25/9) (square root both sides) x = 5/3 (simplify. the sqrt of 25 = 5 and the sqrt of 9 = 3)

OpenStudy (mrnood):

OR x^2= 100/36 =10^2/6^2 x= +/- 10/6 = +/- 5/3

OpenStudy (midhun.madhu1987):

 and you are correct :)